An inheritance of $900000 is to be divided among Ann, Alice, and Tricia in the following manner: Alice is to receive 3/4 of what ann gets, while Tricia get 1/2 of what Ann gets. How much does each receive?
Answer by rfer(16322)

Ann=x
Alice=3/4x
Tricia=1/2x
--------------
x+3/4x+1/2x=900000
2 1/4x=900000
x=900000/(2 1/4)
x=900000*4/9
x=400000
3/4x=300000
1/2x=200000